# A probabilistic proof of Euler’s pentagonal number theorem

### v1

@@ 1. Prologue
This result was first communicated from Euler to Daniel Bernoulli in a letter dated November 30, 1740111Euler’s original letter, unfortunately, has not been preserved but the lines regarding the pentagonal number theorem were alluded in Bernoulli’s letter dated January 28, 1741. See [5, Brief 52].; however, no proof was included therein. The seemingly earliest demonstration offered by Euler [3] was published in 1760, which is a retelling of the idea in his letter to Goldbach [4, Brief 144] on June 9, 1750. Euler’s proof, as briefly presented at the end of this paper, is inductive, relying on an ingenious arithmetic manipulation of the series.
@@ 3. Event spaces
$\boldsymbol{s}_{i}:=\begin{cases}(s_{0},\ldots,s_{k-1},s_{k})&\text{for $i=0$},\\ (s_{0},\ldots s_{i-1},s_{k},s_{i},\ldots,s_{k-1})&\text{for $i\in[1,k-1]$},\\ (s_{k},s_{0},\ldots,s_{k-1})&\text{for $i=k$}.\end{cases}$
@@
Our expected event space is a collection of shuffles denoted by the set
@@
Next, we recall that the union of two sets $S_{1}$ and $S_{2}$ is written as $S_{1}\sqcup S_{2}$ once $S_{1}$ and $S_{2}$ are disjoint; this union is called disjoint union.
@@
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$ with $K$ nonempty. Then
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}=\big{\langle}(I\sqcup K)_{\triangleright},J_{\triangleleft}\big{\rangle}\sqcup\big{\langle}I_{\triangleright},(J\sqcup K)_{\triangleleft}\big{\rangle}.$ (3.4)
@@ 4. Probabilities
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$ with $K$ nonempty. Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{Pr}_{\langle(I\sqcup K)_{\triangleright},J_{\triangleleft}\rangle}+\mathbf{Pr}_{\langle I_{\triangleright},(J\sqcup K)_{\triangleleft}\rangle}.$ (4.4)
@@ 5. The pentagonal number theorem
Singling out the $k=1$ term of the sum in the above gives the desired part $\mathbf{Pr}_{\langle[N+2,2N+3]_{\triangleright},\varnothing_{\triangleleft}\rangle}$. ∎
@@ 6. Epilogue
$\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=q^{\sum_{n=N}^{2N}(k+1+n)}(q^{1+N};q)_{k}$
@@ Acknowledgements
This work was supported by the Austrian Science Fund (No. 10.55776/F1002).

### v2

@@ 1. Prologue
This result was first communicated from Euler to Daniel Bernoulli in a letter dated November 30, 1740111Euler’s original letter, unfortunately, has not been preserved but the lines regarding the pentagonal number theorem were alluded in Bernoulli’s reply dated January 28, 1741. See [5, Brief 52].; however, no proof was included therein. The seemingly earliest demonstration offered by Euler [3] was published in 1760, which is a retelling of the idea in his letter to Goldbach [4, Brief 144] on June 9, 1750. Euler’s proof, as briefly presented at the end of this paper, is inductive, relying on an ingenious arithmetic manipulation of the series.
@@ 3. Event spaces
$\boldsymbol{s}_{i}:=\begin{cases}(s_{0},\ldots,s_{k-1},s_{k})&\text{for $i=0$},\\ (s_{0},\ldots s_{k-i-1},s_{k},s_{k-i},\ldots,s_{k-1})&\text{for $i\in[1,k-1]$},\\ (s_{k},s_{0},\ldots,s_{k-1})&\text{for $i=k$}.\end{cases}$
@@
Our desired event space is a collection of shuffles denoted by the set
@@
Next, we recall that the union of two sets $S_{1}$ and $S_{2}$ is written as $S_{1}\sqcup S_{2}$ once $S_{1}$ and $S_{2}$ are disjoint; this union is called a disjoint union.
@@
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$. Then
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}=\bigsqcup_{K^{\prime}\subset K}\big{\langle}(I\sqcup K^{\prime})_{\triangleright},(J\sqcup(K\backslash K^{\prime}))_{\triangleleft}\big{\rangle}.$ (3.4)
In particular, if $K$ is a singleton, namely, $K=\{k\}$ for a certain $k\in\mathbb{N}_{+}$, then
$\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}=\big{\langle}(I\sqcup\{k\})_{\triangleright},J_{\triangleleft}\big{\rangle}\sqcup\big{\langle}I_{\triangleright},(J\sqcup\{k\})_{\triangleleft}\big{\rangle}.$ (3.5)
@@ 4. Probabilities
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$. Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\sum_{K^{\prime}\subset K}\mathbf{Pr}_{\langle(I\sqcup K^{\prime})_{\triangleright},(J\sqcup(K\backslash K^{\prime}))_{\triangleleft}\rangle}.$ (4.4)
In particular, if $K$ is a singleton, namely, $K=\{k\}$ for a certain $k\in\mathbb{N}_{+}$, then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{Pr}_{\langle(I\sqcup\{k\})_{\triangleright},J_{\triangleleft}\rangle}+\mathbf{Pr}_{\langle I_{\triangleright},(J\sqcup\{k\})_{\triangleleft}\rangle}.$ (4.5)
@@ 5. The pentagonal number theorem
Singling out the $k=1$ term of the sum in the above gives the desired part $\mathbf{Pr}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$. ∎
@@ 6. Epilogue
$\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\prod_{i=k+1+N}^{k+1+2N}q^{i}\prod_{j=1+N}^{k+N}(1-q^{j})$
@@ Acknowledgements
This work was supported by the Austrian Science Fund (No. 10.55776/F1002). I am grateful to Hugh Thomas for the comments on Proposition 4.4 in an earlier draft.